# Persistence-Based Statistics for Detecting Structural Changes in High-Dimensional Point Clouds

### v1

@@ Abstract
We study the probabilistic behavior of persistence statistics under distributional variability and propose a novel nonparametric framework for detecting structural changes in high-dimensional random point clouds. We first establish moment bounds and tightness results for classical persistence statistics—total and maximum persistence—under general distributions, with explicit scaling behavior derived for Gaussian mixture models. Building on these theoretical foundations, we introduce a normalized statistic based on persistence landscapes combined with the Jensen–Shannon divergence, and we prove its Hölder continuity with respect to perturbations of input point clouds. The resulting measure is stable, scale- and shift-invariant, and suitable for nonparametric inference via permutation testing. A numerical illustration using dynamic attribute vectors from decentralized governance data demonstrates how the proposed method can capture regime shifts and evolving geometric complexity. Our results contribute to the theoretical understanding of random persistence and provide a rigorous statistical foundation for topological change-point detection in complex, high-dimensional systems.
@@
Topological data analysis, Persistent homology, Persistence landscapes, Statistical inference, Random point clouds, Change-point detection
@@ 1 Introduction
Topological data analysis (TDA) provides a robust framework for quantifying geometric and topological structures in complex data. In many real-world systems, however, the underlying data-generating process evolves over time, leading to gradual or abrupt changes in the topology of the associated point clouds. Understanding how persistence statistics behave under such variability— and designing rigorous, distribution-free tools to detect structural changes— remains an open challenge at the intersection of computational topology and probability theory.
@@ 3.4 Moment Bounds and Tightness under GMMs
The closed form in Definition 1 follows from the fact that $\|Z\|$ has the $\chi_{d}$ distribution, i.e., the distribution of the square root of a $\chi^{2}_{d}$ random variable. The $p$-th moment of the $\chi_{d}$ distribution is well known; see, e.g., Vershynin [16, Sec. 2.3] or Muirhead [14, Ch. 1], yielding
@@ 4.1 Definition of the PL+JS Statistic
We now introduce a scale–bounded, distribution–free statistic for comparing persistence diagrams. The construction combines a finite–layer persistence landscape with an information–theoretic divergence on a fixed support. No distributional assumptions (e.g., Gaussian mixtures) are required; those models were used earlier only to motivate variance–scaling phenomena.
@@
Remark 6.
@@
Fixing $S$ globally means that the support of all probability densities is the same interval $[0,S]$, so that the Jensen-Shannon distances computed for different pairs of diagrams are strictly comparable. For practical purposes, it is advisable to choose $S$ somewhat larger than all finite death times expected to appear in the data, thereby ensuring that the supports of the persistence landscapes are fully contained within $[0,S]$.
Remark 7 (On the choice of $\gamma$ and $\theta$ (preview)).
The parameters $\gamma\in[0,1],\theta\in(0,1]$ govern the balance between sensitivity and stability of the associated density $p_{\ell}^{(\gamma,\theta)}(D)$. The mixing weight $\gamma$ controls the proportion of the uniform baseline $u_{S}$, while the regularization level $\theta$ stabilizes the normalization even when the total landscape mass is small. Increasing either $\gamma$ or $\theta$ produces smoother, more conservative densities; decreasing them enhances sensitivity to topological signal. Quantitative guidance and theoretical guarantees for the choice of $(\gamma,\theta)$ will be provided after establishing continuity of the statistic (see Remark 10 below).
Remark 8.
Compared with direct diagram-based distances such as the bottleneck or Wasserstein metrics on persistence diagrams (see, e.g., [4]), the PL+JS statistic offers several advantages. First, its computation relies on persistence landscapes and probability density normalization, which avoids the cubic-time optimal matching problem inherent in bottleneck and Wasserstein computations. Second, the JS divergence is bounded in $[0,1]$, yielding a scale-free and directly interpretable measure, whereas diagram distances depend strongly on the absolute scale of the underlying data. Third, the PL+JS statistic admits a natural information-theoretic interpretation, which contrasts with the more geometric but less statistically transparent nature of diagram distances.
Finally, because the PL+JS statistic is defined between probability distributions, it fits directly into the standard framework of nonparametric hypothesis testing. The Jensen-Shannon divergence is finite, symmetric, and bounded, which stabilizes permutation-based approximations of the null distribution and makes the resulting test outcomes more interpretable. This property is particularly advantageous in practical applications, where reliable and communicable test statistics are crucial.
Remark 9.
A distinctive feature of the proposed PL+JS statistic is its invariance under global shifts of the data distribution. For example, if all attribute vectors are translated by a constant vector, the persistence diagrams and their associated landscapes remain unchanged, so the PL+JS statistic evaluates to zero. This behavior contrasts sharply with classical distributional distances such as the Wasserstein distance, the Energy distance, the Maximum Mean Discrepancy (MMD), the Total Variation distance, the Kolmogorov-Smirnov distance, the Kullback-Leibler divergence, or the Jensen-Shannon divergence, all of which are strictly positive under such shifts.
This invariance is advantageous in attribute data analysis: it enables us to focus not on absolute levels of attributes, but on structural properties such as the emergence of clusters or loop-like features. Consequently, the PL+JS statistic captures meaningful topological changes while ignoring trivial global displacements of the data.
See Table 1 for a summary in comparison with other common distances.
Table 1: Comparison of common distributional distances and the PL+JS statistic
Distance / Divergence Shift Invariance Range
Wasserstein distance No $0,\infty)$
Energy distance No $0,\infty)$
Maximum Mean Discrepancy (MMD) No (depends on kernel) $0,\infty)$ or bounded by kernel
Total Variation distance No $[0,1]$
Kolmogorov-Smirnov distance No $[0,1]$
Kullback-Leibler divergence No $0,\infty]$
Jensen-Shannon divergence No $[0,1]$
PL+JS statistic (proposed) Yes $[0,1]$
@@ 4.2 Continuity of the Proposed Change-Point Detection Statistic
In what follows, we reuse the notation introduced in Section 3.1. Let $\mathcal{X}$ be $\mathbb{R}^{d}$ with $d\in\mathbb{N}$, where each element of $\mathcal{X}$ represents an individual attribute vector. We defined the map $\mathcal{D}_{\ell}\colon\mathcal{X}\to\mathrm{Dgm}$ that assigns, to a finite point cloud $x=(x_{1},\dots,x_{N})$, the persistence diagram in degree $\ell\in\mathbb{N}_{0}$ associated with the Vietoris-Rips filtration. Here, since the numbers of points $N_{1}$ and $N_{2}$ may differ for two data windows $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$, we explicitly indicate the dependence on the number of points by writing $\mathcal{D}_{\ell}^{(N_{1})}$ and $\mathcal{D}_{\ell}^{(N_{2})}$.
@@
$T_{\ell}(x^{(1)},x^{(2)})=\Delta_{\ell}(\mathcal{D}_{\ell}^{(N_{1})}(x^{(1)}),\mathcal{D}_{\ell}^{(N_{2})}(x^{(2)})),$
where $\Delta_{\ell}\colon\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R}$ is the map defined in Definition 2. Equip $\mathcal{X}^{N_{i}}$ with the metric
$d_{\mathcal{X}^{N_{i}}}(x^{(i)},\tilde{x}^{(i)}):=\max_{1\leq j\leq N_{i}}\|x_{j}^{(i)}-\tilde{x}_{j}^{(i)}\|,$
and $\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}$ with the metric
$d_{\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}}\left((x^{(1)},x^{(2)}),(\tilde{x}^{(1)},\tilde{x}^{(2)})\right):=d_{\mathcal{X}^{N_{1}}}(x^{(1)},\tilde{x}^{(1)})+d_{\mathcal{X}^{N_{2}}}(x^{(2)},\tilde{x}^{(2)}).$
Then $T_{\ell}$ is $1/2$-Hölder continuous with respect to this metric:
@@
be the metric on $\mathrm{Dgm}\times\mathrm{Dgm}$. The map
@@
For each simplex $\sigma\in K^{(i)}:=2^{\{1,\ldots,N_{i}\}}\setminus\{\emptyset\}$ ($i=1,2$ ) we have
$\Phi^{(i)}_{\sigma}(x^{(i)})=\max_{a,b\in\sigma}\|x_{a}^{(i)}-x_{b}^{(i)}\|.$
Using that the map $x\mapsto\max_{i}x_{i}$ is 1-Lipschitz continuous with respect to $\|\cdot\|_{\infty}$, together with the reverse triangle inequality $|\|u\|-\|v\||\leq\|u-v\|$, we have
[…]
$\leq\max_{a,b\in\sigma}\|(x_{a}^{(i)}-\tilde{x}_{a}^{(i)})-(x_{b}^{(i)}-\tilde{x}_{b}^{(i)})\|$
$\leq 2d_{\mathcal{X}^{N_{i}}}(x^{(i)},\tilde{x}^{(i)}).$
Therefore
$\sup_{\sigma\in K^{(i)}}\left|\Phi^{(i)}_{\sigma}(x^{(i)})-\Phi^{(i)}_{\sigma}(\tilde{x}^{(i)})\right|\leq 2d_{\mathcal{X}^{N_{i}}}(x^{(i)},\tilde{x}^{(i)}).$
By the stability theorem for persistence diagrams (see, e.g., [4]), we have
$d_{B}\bigl(\mathcal{D}_{\ell}^{(N_{i})}(x^{(i)}),\mathcal{D}_{\ell}^{(N_{i})}(\tilde{x}^{(i)})\bigr)\leq\sup_{\sigma\in K^{(i)}}\bigl|\Phi^{(i)}_{\sigma}(x^{(i)})-\Phi^{(i)}_{\sigma}(\tilde{x}^{(i)})\bigr|\leq 2d_{\mathcal{X}^{N_{i}}}(x^{(i)},\tilde{x}^{(i)}).$
Applying Proposition 10 gives
@@
$=\left|\Delta_{\ell}\bigl(\mathcal{D}_{\ell}^{(N_{1})}(x^{(1)}),\mathcal{D}_{\ell}^{(N_{2})}(x^{(2)})\bigr)-\Delta_{\ell}\bigl(\mathcal{D}_{\ell}^{(N_{1})}(\tilde{x}^{(1)}),\mathcal{D}_{\ell}^{(N_{2})}(\tilde{x}^{(2)})\bigr)\right|$
$\leq C_{\Delta}\sqrt{d_{B}\bigl(\mathcal{D}_{\ell}^{(N_{1})}(x^{(1)}),\mathcal{D}_{\ell}^{(N_{1})}(\tilde{x}^{(1)})\bigr)+d_{B}\bigl(\mathcal{D}_{\ell}^{(N_{2})}(x^{(2)}),\mathcal{D}_{\ell}^{(N_{2})}(\tilde{x}^{(2)})\bigr)}$
$\leq\sqrt{2}C_{\Delta}\sqrt{d_{\mathcal{X}^{N_{1}}}(x^{(1)},\tilde{x}^{(1)})+d_{\mathcal{X}^{N_{2}}}(x^{(2)},\tilde{x}^{(2)})}.$

### v2

@@ Abstract
We study the probabilistic behavior of persistence-based statistics and propose a novel nonparametric framework for detecting structural changes in high-dimensional random point clouds. We establish moment bounds and tightness results for classical persistence statistics—total and maximum persistence—under general distributions, with explicit variance-scaling behavior derived for Gaussian mixture models. Building on these results, we introduce a bounded and normalized statistic based on persistence landscapes combined with the Jensen–Shannon divergence, and we prove its Hölder continuity with respect to perturbations of the input point clouds. The resulting measure is stable, scale- and shift-invariant, and well suited for finite-sample nonparametric inference via permutation testing. An illustrative numerical study using dynamic attribute vectors from decentralized governance data demonstrates the practical applicability of the proposed method. Overall, this work provides a statistically rigorous and computationally stable approach to change-point detection in complex, high-dimensional data.
@@
Nonparametric inference, Change-point detection, Random point clouds, Topological data analysis, Persistent homology, Persistence landscapes
@@ 1 Introduction
Statistical detection of structural changes in high-dimensional data is a fundamental problem in modern data analysis, with applications ranging from complex networks to time-evolving multivariate systems. When data are represented as point clouds, such changes often manifest as alterations in their geometric or topological organization. Designing nonparametric and computationally stable procedures that can capture such changes in a distribution-free manner remains a challenging task.
Topological data analysis (TDA) provides a powerful framework for quantifying geometric and topological structure in point cloud data. In many real-world systems, however, the underlying data-generating process evolves over time, leading to gradual or abrupt changes in the associated topology. Understanding how persistence statistics behave under such variability —and how they can be used for statistically rigorous inference—remains an open problem at the interface of statistics, probability theory, and computational topology.
@@ 2.2 Topological Data Analysis and Related Work
From a statistical perspective, many existing persistence-based approaches focus primarily on feature extraction or empirical summaries, while questions of finite-sample behavior, distribution-free inference, and computational stability under data perturbations remain comparatively less explored.
@@ 3 Persistence Statistics and Moment Bounds
This section adopts a standard formal framework for the persistent homology of finite point clouds, where persistence diagrams are treated as diagram-valued mappings induced from the data through filtration constructions. The overall setup—including the representation of a point cloud as a finite subset of a metric space and the use of persistence-based real-valued quantities derived from filtrations—follows the general formulation introduced by [4]. Within this framework, we focus on probabilistic moment bounds and scaling properties of persistence statistics.
@@ 3.4 Moment Bounds and Tightness under GMMs
The closed form in Definition 1 follows from the fact that $\|Z\|$ has the $\chi_{d}$ distribution, i.e., the distribution of the square root of a $\chi^{2}_{d}$ random variable. The $p$-th moment of the $\chi_{d}$ distribution is well known; see, e.g., Vershynin [19, Sec. 2.3] or Muirhead [17, Ch. 1], yielding
@@
Remark 6 (Role of moment bounds, tightness, and variance scaling).
The results of this section serve a conceptual and methodological role that goes beyond the specific Gaussian mixture setting in which they are derived.
(i) Beyond trivial global rescaling. While a deterministic rescaling of a point cloud trivially rescales all Vietoris–Rips filtration values, the setting considered in Theorem 5 is fundamentally different. Here, the mean structure of the distribution is fixed, and only the covariance matrices are scaled. Thus, the observed $\sqrt{\eta}$–growth of persistence statistics reflects increasing stochastic dispersion around fixed reference locations, rather than a uniform geometric dilation. This distinction is essential in applications where scale changes arise from heterogeneity or volatility, rather than from a deterministic rescaling of the data.
(ii) Tightness and uniform tail control. Corollary 6 and Corollary 8 establish that, after normalization by $\sqrt{\eta}$, classical persistence statistics form a tight family with uniform polynomial tail bounds. These results provide non-asymptotic control over the magnitude and variability of persistence-based quantities under increasing dispersion. In particular, they justify the use of scale-normalized or bounded statistics when comparing random point clouds whose intrinsic variability may differ.
(iii) Motivation for bounded and normalized statistics. The variance-scaling behavior quantified in Theorem 5 highlights a limitation of raw persistence summaries: their magnitude is dominated by dispersion effects. This observation motivates the construction in Section 4.1 of the PL+JS statistic, which suppresses scale-driven growth by normalization and regularization, while retaining sensitivity to changes in topological structure. In this sense, the results of this section provide the probabilistic foundation for why a bounded, scale-insensitive comparison of persistence diagrams is desirable for nonparametric inference and change-point detection.
Taken together, the moment bounds, tightness, and uniform tail estimates clarify the stochastic behavior of classical persistence statistics and explain their role as a theoretical benchmark. They do not directly enter the definition of the PL+JS statistic, but they motivate its design and delineate the regime in which purely scale-driven variability should be factored out.
@@ 4.1 Definition of the PL+JS Statistic
We now introduce a scale–bounded, distribution–free statistic for comparing persistence diagrams. The motivation for this construction builds on the results of the preceding sections.
In Sections 3.3 and 3.4, we showed that classical persistence statistics, such as total and maximum persistence, typically increase with the dispersion of the underlying data distribution, and in particular scale with the variance multiplier under Gaussian mixture models. While this behavior is natural from a geometric viewpoint, it implies that raw persistence quantities are inherently sensitive to global scale and may obscure more subtle structural changes.
The PL+JS statistic is designed to mitigate this effect by normalizing persistence landscapes and comparing them through a bounded, information–theoretic divergence on a fixed support. In this way, the resulting statistic suppresses purely scale–driven variation and enables meaningful comparison of topological structure across samples, making it suitable for nonparametric change–point detection. No distributional assumptions (e.g., Gaussian mixtures) are required at this stage; those models were used earlier only to motivate variance–scaling phenomena.
@@
Remark 7 (Interpretability, boundedness, and parameter control).
The use of the Jensen–Shannon distance in Definition 2 yields a statistic $\Delta_{\ell}$ that is bounded in $[0,1]$, providing a scale-free and directly interpretable measure of discrepancy between persistence diagrams. This boundedness is particularly advantageous for finite-sample nonparametric inference, including permutation-based testing, as it prevents a small number of extreme persistence features from dominating the statistic.
The regularization and mixing parameters $\theta$ and $\gamma$ stabilize the normalization of persistence landscapes when the total landscape mass is small. Their influence is explicit and quantitative: as shown in Proposition 10 and Theorem 9, increasing either parameter improves robustness by reducing the Hölder constant, while smaller values enhance sensitivity to topological signal. Thus, $(\gamma,\theta)$ provide principled control of the stability–sensitivity trade-off rather than ad hoc regularization.
Remark 8 (Landscape sums versus silhouettes).
Persistence silhouettes aggregate all triangular functions associated with a diagram into a single weighted average, emphasizing long-lived features through nonlinear persistence weights [5]. In contrast, the present work employs the unweighted sum of the first $M$ persistence landscape layers, $\Lambda_{\ell}^{(M)}=\sum_{m=1}^{M}\lambda_{\ell,m}$.
This choice is motivated by transparency, stability, and practical controllability. The parameter $M$ governs the amount of topological information included in the statistic in a monotone and easily interpretable manner, and the resulting functional inherits a Lipschitz-type stability bound that grows linearly in $M$.
From a practical perspective, $M$ can be selected in an adaptive manner. One may begin with a small value of $M$, capturing only the most dominant topological features, and assess whether structural changes are detected. If no significant differences are observed, $M$ can be increased to incorporate finer-scale features. Since increasing $M$ leads to higher computational cost, its choice naturally reflects a trade-off between statistical sensitivity and computational efficiency.
While silhouettes provide an effective low-dimensional summary, the landscape-sum representation retains an explicit hierarchical ordering of dominant features. This hierarchical structure, together with the tunable parameter $M$, is particularly advantageous for detecting structural changes in a statistically controlled and computationally feasible manner.
Remark 9 (Global cutoff and comparability).
@@
Fixing the cutoff $S$ globally ensures that all probability densities associated with persistence diagrams share the same support $[0,S]$. As a result, Jensen–Shannon distances computed across different pairs of diagrams are strictly comparable. In practice, $S$ should be chosen slightly larger than the maximal finite death time expected in the data to avoid truncation effects.
Remark 10 (Limits of signature-based testing).
Any test based on a compressed signature of persistence diagrams necessarily identifies distributions only up to the equivalence relation induced by that signature. Distinct persistence diagrams may therefore become indistinguishable after compression; this phenomenon is intrinsic to any signature-based representation and is well illustrated, for example, by the notion of Euler equivalence discussed by Dłotko et al. [7] for the Euler characteristic curve.
Accordingly, the PL+JS statistic is not intended as a complete invariant of persistence diagrams, but as a stable and inference-oriented summary. It is designed to provide reliable power against a broad class of structured alternatives while sacrificing sensitivity to fine-scale differences that are often dominated by noise or sampling variability.
@@ 4.2 Continuity of the Proposed Change-Point Detection Statistic
In what follows, we reuse the notation introduced in Section 3.1. Let $\mathcal{X}=\mathbb{R}^{d}$ with $d\in\mathbb{N}$, where each element of $\mathcal{X}$ represents an individual attribute vector. For each $N\in\mathbb{N}$, let
$\mathcal{D}_{\ell}^{(N)}\colon\mathcal{X}^{N}\to\mathrm{Dgm}$
denote the map that assigns to a finite point cloud $x=(x_{1},\dots,x_{N})$ the persistence diagram in degree $\ell\in\mathbb{N}_{0}$ associated with the Vietoris–Rips filtration. When no confusion arises, we simply write $\mathcal{D}_{\ell}$ instead of $\mathcal{D}_{\ell}^{(N)}$.
@@
$T_{\ell}(x^{(1)},x^{(2)})=\Delta_{\ell}(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(x^{(2)})),$
where $\Delta_{\ell}\colon\mathrm{Dgm}\times\mathrm{Dgm}\to\mathbb{R}$ is the map defined in Definition 2. Equip $\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}$ with the metric
$d_{\mathcal{X}^{N_{1}}\times\mathcal{X}^{N_{2}}}\left((x^{(1)},x^{(2)}),(\tilde{x}^{(1)},\tilde{x}^{(2)})\right):=d_{H}^{(N_{1})}(x^{(1)},\tilde{x}^{(1)})+d_{H}^{(N_{2})}(x^{(2)},\tilde{x}^{(2)}),$
where
$d_{H}^{(N)}(x,\tilde{x}):=\max\Bigl\{\max_{1\leq i\leq N}\min_{1\leq j\leq N}\|x_{i}-\tilde{x}_{j}\|,\ \max_{1\leq j\leq N}\min_{1\leq i\leq N}\|x_{i}-\tilde{x}_{j}\|\Bigr\}$
defines a Hausdorff-type metric on $\mathcal{X}^{N}$. Then $T_{\ell}$ is $1/2$-Hölder continuous with respect to this metric:
@@
be the metric on $\mathrm{Dgm}\times\mathrm{Dgm}$ where $d_{B}$ denotes the bottleneck distance (Appendix A.4). The map
@@
Before proving Theorem 9, we prepare a standard stability result for Vietoris–Rips persistence diagrams with respect to the Hausdorff distance. This result will serve as a key technical ingredient in the proof.
Proposition 14 (Stability of Vietoris–Rips persistence diagrams).
Let $X,Y\subset\mathbb{R}^{d}$ be nonempty finite sets, and let $\mathrm{VR}(X)=\{\mathrm{VR}(X,r)\}_{r\geq 0}$ and $\mathrm{VR}(Y)=\{\mathrm{VR}(Y,r)\}_{r\geq 0}$ denote the Vietoris–Rips filtrations parametrized by the diameter, i.e.,
$\mathrm{VR}(X,r):=\{\sigma\subset X\mid\max_{x,x^{\prime}\in\sigma}\|x-x^{\prime}\|\leq r\}.$
Then, for each homological degree $\ell\geq 0$, the corresponding persistence diagrams satisfy
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\;\leq\;2\,d_{H}(X,Y),$
where $d_{B}$ denotes the bottleneck distance and $d_{H}$ the Hausdorff distance.
Proof.
First, set $\varepsilon:=d_{H}(X,Y)$.
By the definition of the Hausdorff distance, for any $x\in X$ there exists a point $y\in Y$ such that $\|x-y\|\leq\varepsilon$. For each $x\in X$, choose one such point and define $f(x)\in Y$. Similarly, for each $y\in Y$, choose a point $x\in X$ satisfying $\|y-x\|\leq\varepsilon$ and define $g(y)\in X$. Then it follows that $\|x-f(x)\|\leq\varepsilon$ for all $x\in X$, and $\|y-g(y)\|\leq\varepsilon$ for all $y\in Y$.
Fix an arbitrary $r\geq 0$.
Take an arbitrary simplex $\{x_{0},\dots,x_{k}\}\subset X$ contained in $\mathrm{VR}(X,r)$. For any $i,j$, by the definition of $\mathrm{VR}(X,r)$ we have $\|x_{i}-x_{j}\|\leq r$, and hence by the triangle inequality,
$\|f(x_{i})-f(x_{j})\|\leq\|f(x_{i})-x_{i}\|+\|x_{i}-x_{j}\|+\|x_{j}-f(x_{j})\|\leq r+2\varepsilon.$
Therefore, $\{f(x_{0}),\ldots,f(x_{k})\}$ is a simplex in $\mathrm{VR}(Y,r+2\varepsilon)$. This correspondence defines a simplicial map
$F_{r}:\mathrm{VR}(X,r)\longrightarrow\mathrm{VR}(Y,r+2\varepsilon),$
and similarly a simplicial map
$G_{r}:\mathrm{VR}(Y,r)\longrightarrow\mathrm{VR}(X,r+2\varepsilon).$
Let $r^{\prime}\geq r$ be arbitrary. Note that for each simplex $\sigma=\{x_{0},\ldots,x_{k}\}\in\mathrm{VR}(X,r)$, we define
$F_{r}(\sigma):=\{f(x_{0}),\ldots,f(x_{k})\}.$
Since this definition does not depend on $r$, for $\sigma\in\mathrm{VR}(X,r^{\prime})$ we also have
$F_{r^{\prime}}(\sigma)=\{f(x_{0}),\ldots,f(x_{k})\}.$
Consequently,
$F_{r^{\prime}}\bigl(\iota_{r^{\prime},r}^{(X)}(\sigma)\bigr)=F_{r^{\prime}}(\sigma)=\{f(x_{0}),\ldots,f(x_{k})\}.$
On the other hand,
$\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}\bigl(F_{r}(\sigma)\bigr)=\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}\bigl(\{f(x_{0}),\ldots,f(x_{k})\}\bigr)=\{f(x_{0}),\ldots,f(x_{k})\}.$
Hence, the following diagram commutes for the family $\{F_{r}\}_{r\geq 0}$:
$\mathrm{VR}(X,r)$$\mathrm{VR}(X,r^{\prime})$$\mathrm{VR}(Y,r+2\varepsilon)$$\mathrm{VR}(Y,r^{\prime}+2\varepsilon).$$\scriptstyle{\iota_{r^{\prime},r}^{(X)}}$$\scriptstyle{F_{r}}$$\scriptstyle{F_{r^{\prime}}}$$\scriptstyle{\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(Y)}}$
By the same argument, since the definition of $G_{r}$ does not depend on the parameter $r$, the family $\{G_{r}\}_{r\geq 0}$ also satisfies the corresponding commutativity condition. More precisely, for any $r^{\prime}\geq r$, the diagram
$\mathrm{VR}(Y,r)$$\mathrm{VR}(Y,r^{\prime})$$\mathrm{VR}(X,r+2\varepsilon)$$\mathrm{VR}(X,r^{\prime}+2\varepsilon)$$\scriptstyle{\iota_{r^{\prime},r}^{(Y)}}$$\scriptstyle{G_{r}}$$\scriptstyle{G_{r^{\prime}}}$$\scriptstyle{\iota_{r^{\prime}+2\varepsilon,r+2\varepsilon}^{(X)}}$
commutes.
Next, for any simplex $\sigma=\{x_{0},\ldots,x_{k}\}\in\mathrm{VR}(X,r)$, the equality
$\iota_{r+4\varepsilon,r}^{(X)}(\sigma)\cup(G_{r+2\varepsilon}\circ F_{r})(\sigma)=\sigma\cup(G_{r+2\varepsilon}\circ F_{r})(\sigma)$
defines a simplex in $\mathrm{VR}(X,r+4\varepsilon)$. Indeed, this follows from the estimates
$\|x_{i}-g(f(x_{i}))\|\leq\|x_{i}-f(x_{i})\|+\|f(x_{i})-g(f(x_{i}))\|\leq 2\varepsilon,$
$\|x_{i}-g(f(x_{j}))\|\leq\|x_{i}-x_{j}\|+\|x_{j}-f(x_{j})\|+\|f(x_{j})-g(f(x_{j}))\|$
$\quad\leq r+\varepsilon+\varepsilon=r+2\varepsilon,$
$\|g(f(x_{i}))-g(f(x_{j}))\|\leq\|g(f(x_{i}))-f(x_{i})\|$
$\quad+\|f(x_{i})-f(x_{j})\|+\|f(x_{j})-g(f(x_{j}))\|$
$\quad\leq\varepsilon+(r+2\varepsilon)+\varepsilon=r+4\varepsilon.$
Therefore, the composite map
$G_{r+2\varepsilon}\circ F_{r}\colon\mathrm{VR}(X,r)\to\mathrm{VR}(X,r+4\varepsilon)$
is contiguous to the inclusion map $\iota_{r+4\varepsilon,r}^{(X)}$ in $\mathrm{VR}(X,r+4\varepsilon)$. Similarly, the composite map
$F_{r+2\varepsilon}\circ G_{r}\colon\mathrm{VR}(Y,r)\to\mathrm{VR}(Y,r+4\varepsilon)$
is contiguous to the inclusion $\iota_{r+4\varepsilon,r}^{(Y)}$ in $\mathrm{VR}(Y,r+4\varepsilon)$.
In general, two contiguous simplicial maps are homotopic on geometric realizations and therefore induce identical maps on homology [8]. Consequently, the homomorphism induced by $G_{r+2\varepsilon}\circ F_{r}$,
$(G_{r+2\varepsilon}\circ F_{r})_{*}\colon H_{\ell}(\mathrm{VR}(X,r))\to H_{\ell}(\mathrm{VR}(X,r+4\varepsilon)),$
coincides with
$(\iota_{r+4\varepsilon,r}^{(X)})_{*}\colon H_{\ell}(\mathrm{VR}(X,r))\to H_{\ell}(\mathrm{VR}(X,r+4\varepsilon)).$
Similarly, the homomorphism induced by $F_{r+2\varepsilon}\circ G_{r}$,
$(F_{r+2\varepsilon}\circ G_{r})_{*}\colon H_{\ell}(\mathrm{VR}(Y,r))\to H_{\ell}(\mathrm{VR}(Y,r+4\varepsilon)),$
coincides with
$(\iota_{r+4\varepsilon,r}^{(Y)})_{*}\colon H_{\ell}(\mathrm{VR}(Y,r))\to H_{\ell}(\mathrm{VR}(Y,r+4\varepsilon)).$
Hence, the persistence modules $\mathbb{X}:=\{H_{\ell}(\mathrm{VR}(X,r))\}_{r\geq 0}$ and $\mathbb{Y}:=\{H_{\ell}(\mathrm{VR}(Y,r))\}_{r\geq 0}$ are $2\varepsilon$-interleaved, and their interleaving distance is at most $2\varepsilon$.
Since Vietoris–Rips filtrations on finite point sets are finitely generated in each degree, the persistence modules $\mathbb{X}$ and $\mathbb{Y}$ are q-tame. Therefore, by the stability theorem of persistence diagrams [6], the bottleneck distance $d_{B}$ and the interleaving distance $d_{I}$ satisfy
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\leq d_{I}(\mathbb{X},\mathbb{Y}),$
and hence
$d_{B}\bigl(D_{\ell}(\mathrm{VR}(X)),D_{\ell}(\mathrm{VR}(Y))\bigr)\leq 2\varepsilon=2\,d_{H}(X,Y).$
This completes the proof. ∎
Remark 11.
The factor $2$ arises from the diameter parametrization of the Vietoris–Rips filtration. For the Čech filtration or for a radius-normalized Vietoris–Rips filtration, the corresponding stability bound holds with constant $1$.
@@
Applying Proposition 10 and Proposition 14 gives
@@
$=\left|\Delta_{\ell}\bigl(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(x^{(2)})\bigr)-\Delta_{\ell}\bigl(\mathcal{D}_{\ell}(\tilde{x}^{(1)}),\mathcal{D}_{\ell}(\tilde{x}^{(2)})\bigr)\right|$
$\leq C_{\Delta}\sqrt{d_{B}\bigl(\mathcal{D}_{\ell}(x^{(1)}),\mathcal{D}_{\ell}(\tilde{x}^{(1)})\bigr)+d_{B}\bigl(\mathcal{D}_{\ell}(x^{(2)}),\mathcal{D}_{\ell}(\tilde{x}^{(2)})\bigr)}$
$\leq\sqrt{2}C_{\Delta}\sqrt{d_{H}^{(N_{1})}(x^{(1)},\tilde{x}^{(1)})+d_{H}^{(N_{2})}(x^{(2)},\tilde{x}^{(2)})}.$